Update at 14:27:28(BST): The network has functioned correctly since the last update so this issue is now being marked as resolved.
Update at 08:47:01(BST): The engineers have now found the issue and put a fix in place. They are still monitoring this to ensure all is running ok still.
08:02:06(BST): Unfortunately the issue is still ongoing, the engineers are still working on this. We will update this page with more information as soon as we receive this from the engineers.
There is currently a high load of traffic inside our Manchester data centre causing sites to occasionally fail to load or load slowly.
Our engineers have been alerted and are investigating into the matter so try and isolate the source of this traffic so action can be taken to bring services back to normal.
Please accept our apologies for any inconvenience caused and our assurances we are doing all we can to get this resolved as quickly as possible.
1st May 2012, at 6:21AM
1st May 2012, at 1:27PM
Update at 20:38:02(BST): Update at 11:40:06(BST): The engineers think they have tracked down the issue to a particular site that was causing a high load on the server. We are looking into this further and hope to have this resolved today.
Update at 14:31:46(BST): We are aware this issue reoccurred within the last 10 minutes. Our engineers are attempting to trace the issue to stop this happening again. We'll update this post as soon as we have further information.
Update at 08:15:51(BST): The issue has now been resolved with this server.The engineers have performed a reboot of the server and normal service has now resumed on the server.
Unfortunately the server is currently experiencing some issues causing both website and email services to be non responsive.
Our engineers are currently investigating into the matter and are hoping to have all services back up and running soon.
Please accept our apologies for any inconvenience and our assurances we are doing all that is possible to get this resolved as soon as possible.
27th April 2012, at 7:58AM
7th May 2012, at 7:38PM
Update at 09:33:33(BST): The issue has now been resolved with this server.
All service are back on line. Thank you for your patience.
Update at 09:16:04(BST):
Engineers continue to work on the server. We thank you for your patience.
Engineers are currently rebooting this server as an emergency measure. Apologies for any inconvenience caused.
26th April 2012, at 8:57AM
26th April 2012, at 8:33AM
Update at 20:09:04(BST): The issue has now been resolved with this server.This issue was fixed at 8.03PM
Update at 19:52:58(BST): An issue has resurfaced on this server and we are looking at it again.
Update at 13:04:39(BST): The issue has now been resolved with this server.
All sites are now back on line. Apologies for any inconvenience caused.
Update at 12:58:38(BST):
We have identified the issue and hope to have websites back on line as soon as possible. Thank you for your patience.
Engineers are currently looking into the current issues. Apologies for any inconvenience caused.
25th April 2012, at 12:29PM
25th April 2012, at 7:09PM
Update at 12:29:43(BST): Update at 13:39:00(BST): Unfortunately the server appears to be having issues causing websites to fail at times. The engineers are continuing to investigate into this and will do so until the matter is resolved.
Update at 12:04:57(BST): All services are now back online. Apologies for any inconvenience caused.
Update at 11:52:43(BST):
Engineers are now repairing the database on the server. We thank you for you continued patience.
Update at 09:46:56(BST):
Engineers have now rebooted the server again, having carried out various tests. We hope to have the web service back online as soon as possible.
Thank you for your continued patience.
Update at 09:05:34(BST):
Engineers are trying to identify the cause of the issues. We thank you for your patience.
The server stopped responding and engineers have had to reboot it. The server is now scanning its disks for errors. We hope to have web hosting service back on line as soon as possible.
24th April 2012, at 8:42AM
25th April 2012, at 11:29AM
Update at 11:52:53(BST): Update at 10:25:21(BST): We had a temporary issue on our email login system.
If you close down your mail client and re start it, you should be able to login to mail again. Apologies for any inconvenience caused.
23rd April 2012, at 10:23AM
24th April 2012, at 10:52AM
Update at 10:19:06(BST): Update at 09:13:30(BST): The engineers tracked the issue down to UDP flooding from some of our reseller accounts. This has now been stopped and we're monitoring the situation.
We have noticed a degradation in network performance. Our engineers are tracking down the issue to resolve this as soon as possible. We apologise for the inconvenience caused.
20th April 2012, at 9:06AM
23rd April 2012, at 9:19AM
Update at 18:23:58(BST): Update at 18:17:01(BST): The server is now responding following reboot.
This server has become unresponsive and is being rebooted at this time.
Services are expected back online momentarily.
18th April 2012, at 6:11PM
18th April 2012, at 5:23PM
Update at 10:04:39(BST): The issue has now been resolved with this server.
17th April 2012, at 9:05AM
17th April 2012, at 9:04AM
Update at 09:34:22(BST): The issue with this server is now resolved
14th April 2012, at 10:42AM
17th April 2012, at 8:34AM